OpenStudy (anonymous):

if 1 and 3 consecutive integers are added, the result is 69 less than 5 times the second integers. How would you write?

OpenStudy (barrycarter):

I'm assuming the "1 and" is a typo. Let n be the first integer. The next two are n+1 and n+2, so n + (n+1) + (n+2) = 5*(n+1)-69 The only value of n that works is n=67/2, so there is no integer solution.